OpenStudy (aykayyy):

Use the method of cylindrical shells to find the volume gen- erated by rotating the region bounded by the given curves about the specified axis. Sketch the region and a typical shell. y = 4x + x^2 ; y = 3 ; about x = 1

zepdrix (zepdrix):

`Sketch the region and a typical shell.` Let's do this part first. Drawing it out might make it easier to understand how to setup the integral.

OpenStudy (aykayyy):

I did.. and so far I have gotten V = integral(x=1 to x=3) of [2pi(x)(4x-x^2-3)] idk if im just doing the math wrong or i set it up wrong...but i cant get it...the answer is 8pi/3

zepdrix (zepdrix):

Oh boy I can't figure out what's going on with your numbers here :( Hmm

zepdrix (zepdrix):

https://www.desmos.com/calculator/vhjceqdaxb Here is a graph of our region. We'll be integrating with respect to x from one point of intersection to the other. The points are clearly not x=1 and x=3. Where did those come from? <:o

zepdrix (zepdrix):

Sorry that link didn't paste correctly the first time :)

OpenStudy (aykayyy):

sorry.... its actually y = 4x - x^2

OpenStudy (aykayyy):

it would be this now https://www.desmos.com/calculator/vhjceqdaxb

zepdrix (zepdrix):

Oh.........

OpenStudy (aykayyy):

yeah.... so i set it up like this \[V = \int\limits_{1}^{3} 2\pi x (4x-x ^{2}-3)dx\]

zepdrix (zepdrix):

Volume of a single shell would be, V=(Circumference)(height)(Thickness) You have the thickness correct, dx. You have the height correct, upper function minus lower. Your radius for the circumference is wrong. The radius is x-1, since we're spinning around x=1. That might fix things up for you :)

zepdrix (zepdrix):

\[\large V = \int\limits\limits_{1}^{3} 2\pi (x-1) (4x-x ^{2}-3)dx\]

OpenStudy (aykayyy):

hahah i tried that as well.... and i got 89pi/3 hopefully i just did the algebra wrong... ill try again..

OpenStudy (aykayyy):

thanks for your help!

zepdrix (zepdrix):

imma give it a run, see what i come up with :)

OpenStudy (aykayyy):

thanks :)

zepdrix (zepdrix):

I hate working these out, it's so easy to make a mistake after you set it up :P

OpenStudy (aykayyy):

oh wow... i just did it and got 8pi/3 it was a stupid algebra error... im just really tired... thanks anyways!

zepdrix (zepdrix):

cool, good job c:

OpenStudy (aykayyy):

the algebra always kills me.... i put a plus instead of a minus or something all the time haha

zepdrix (zepdrix):

heh